Can someone please step me through the following question? A radioactive material decays according to the formula A = A0(2)-t. How much of a 8 gram sample will be left in 6 year
Answer by stanbon(75887)

A radioactive material decays according to the formula A = Ao(2)^-t. How much of a 8 gram sample will be left in 6 year.
----------------------------------------------
A(6) = 8*2^-6
---
A(6) = 8/64
---
A(6) = 1/8 gram
=====================
Cheers,
Stan H.